OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of seventy-two feet?

OpenStudy (anonymous):

I love to post random questions for no reasons :D

OpenStudy (johnweldon1993):

We know that circumference is \[C = \pi d\] where 'd' is diameter....so since we have C we can do \[d = \frac{ C }{ \pi }\] Just plug in C Lol its fun to see how many people help haha :D

OpenStudy (anonymous):

Thanks for the feedback~